OpenStudy (anonymous):

Find x in the following right triangle, where a = 4. (Round your answer to one decimal place.)

OpenStudy (anonymous):

Picture:

hartnn (hartnn):

same pythagoras theorem

hartnn (hartnn):

\(x^2 = a^2+a^2\)

hartnn (hartnn):

\(x^2= 4^2+4^2\)

OpenStudy (anonymous):

32

hartnn (hartnn):

thats x^2 not x

OpenStudy (anonymous):

\[x=\sqrt{32}\]

OpenStudy (anonymous):

5.6, but if I round it, it's 5.7 right?

hartnn (hartnn):

5.656 yes, it'd be 5.7 :)

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (factor):

dragon ballz kinda is lame

OpenStudy (factor):

infact it is lame not kinda lame

OpenStudy (anonymous):

-.-- get tf on my post disrespecting DBZ

OpenStudy (factor):

its aint ur website and ur nothe tfing creator of dragon balz

OpenStudy (anonymous):

but it's my post & used to be one of my fav cartoons, so BYEEEEE !

hartnn (hartnn):

knock it off, both of you

OpenStudy (factor):

just bcuz its ur fav cartoon thts y u told bye lol rofl

OpenStudy (factor):

okay @hartnn just kinda got bored no offence cook